OpenStudy (anonymous):

peanut the cat is running along at 4.3m/s when he sees a door 3.3m ahead. he immediately tries to stop, but skids along the floor, slowing down at 2.9m/s^2 . a). sketch a graph of peanut's acceleration vs.time b).does he stop or does he slam into the door ? c)if he does stop, how far from the door is he? d). if he slams into the door, how fast is he going when he hits? use kinematics formulas

OpenStudy (cwrw238):

b) u = 4.3, a = -2.9, v = 0, s = distance where u = initial velocity , v = final velocity, a = deceleration use one of the equations for constant acceleration:- v^2 = u^2 + 2as 0 = 4.3^2 - 2*2.9 * s s = 4.3^2/ 5.8 = 3.19 m so does he stop os slam into the door?

OpenStudy (anonymous):

thank you!

OpenStudy (cwrw238):

yw